OpenStudy (anonymous):

Solve the following equation: 9x -3.2 = 6.8 + 8x

OpenStudy (gorv):

9x-8x=6.8+3.2 x=10

OpenStudy (anonymous):

CARP!!!!!!!!!! Evry1 beating me!! XD